OpenStudy (anonymous):

The product of two numbers is 1, and the difference of their squares is 3.75. Find the two numbers.

OpenStudy (anonymous):

\[xy=1\iff y=\frac{1}{x}\]

OpenStudy (anonymous):

solve \[x^2-\frac{1}{x^2}=3.75\]

OpenStudy (anonymous):

The system is : \[xy=1\\x^2-y^2=3.75\] From the 1st equation we can write : \[y=\frac1x\] So, the 2nd becomes : \[x^2-\frac1{x^2}=\frac{15}4\] Then : \[4x^4-4-15x^2=0\] Now let \[t=x^2\] the last equation becomes : \[4t^2-15t-4=0\] We have : \[\Delta= 15^2-4\times 1\times(-4)=241\] So we have : \[t_1=\frac{15+\sqrt{241}}{8}\\ t_2=\frac{15-\sqrt{241}}{8}<0 \] So we have : \[x=\pm\sqrt{t_1}=\pm\sqrt{\frac{15+\sqrt{241}}{8}}\] And the values of y are very clear !

OpenStudy (anonymous):

thanks to the both of you :3